artlover03:

What is the point of intersection? *will post* a- (0,1) b- (1,0) c- (1, 1/2) d- (2,4)

artlover03:

\[f(x)= 2^x \] and \[g(x)=(\frac{ 1 }{ 2 })^x\]

Tyrion:

wat do u think the answer is

artlover03:

im not too sure; im not sure how to solve for it either

Tyrion:

@HaydenwH

HaydenwH:

i mean i canget you the answer but im not much for solving it either, we always use calcs for tis stuff

HaydenwH:

@snowflake0531

snowflake0531:

Btw, an easy way to do this is by graphing the two equations-

snowflake0531:

1 attachment
snowflake0531:

so where does it intersect?

artlover03:

1

snowflake0531:

'1' is not a coordinate point-

artlover03:

OOOO... 0,1

snowflake0531:

so which letter is it

artlover03:

a!!

snowflake0531:

yepp~

artlover03:

Thank you for your help :)

snowflake0531:

yw~